OpenStudy (anonymous):

Pretty hard Algebra!! Prove that:

OpenStudy (anonymous):

\[\frac{ 1 }{30 }k(k+1)(2k+1)(3k^2+3k-1+(k+1)^4 = \frac{ 1 }{ 30 }(k+1)[(k+1)+1][2(k+1)+1][3(k+1)^2+3(k+1)-1]\]

OpenStudy (anonymous):

We have to simplify L.H.S

OpenStudy (anonymous):

@koli123able correct the closing parenthesis

OpenStudy (anonymous):

This L.H.S\[\frac{ 1 }{30 }k(k+1)(2k+1)(3k^2+3k-1)+(k+1)^4 \]

OpenStudy (anonymous):

and make it \[\frac{ 1 }{ 30 }(k+1)[(k+1)+1][2(k+1)+1][3(k+1)^2+3(k+1)-1]\]

OpenStudy (anonymous):

i can't do it

OpenStudy (anonymous):

substitute k=k+1 into LHS

OpenStudy (anonymous):

@koli123able i think u r doing some mathematical Induction problem

OpenStudy (anonymous):

actually no , that wudnt work sorry

OpenStudy (anonymous):

yes this topic is related to mathematical induction

OpenStudy (anonymous):

true.... this looks like proof by induction

OpenStudy (anonymous):

i was just asking the thing where i got stuck

OpenStudy (anonymous):

The real question is 1^4+2^4+3^4+....+n^4= 1/30 n(n+1) (2n+1)(3n^2+3n-1)

OpenStudy (anonymous):

prove by mathematical induction

OpenStudy (anonymous):

always give us the real question! haha

OpenStudy (anonymous):

(1) Put n=1 in P(n) 1^4=1/30.1(1+1)(2+1)3+3-1) 1=1 P(n) is true for n=1

OpenStudy (anonymous):

firs take 1/30 *(k+1) common out

OpenStudy (anonymous):

(2)Assume that P(n) is true for n=k

OpenStudy (anonymous):

1^4+2^4+3^4+....+k^4=1/30 and so on

OpenStudy (anonymous):

Add next term on both sides i.e, (k+1)^4

OpenStudy (anonymous):

\[\frac{ 1 }{30 }k(k+1)(2k+1)(3k^2+3k-1)+(k+1)^4 \]

OpenStudy (anonymous):

so this was the whole thing

OpenStudy (anonymous):

it all comes down to the prove question i asked

OpenStudy (anonymous):

i cant prove it

OpenStudy (anonymous):

i have done all of the questions by same method except this one.....

OpenStudy (anonymous):

u have alot of expnding to do and u have to use synthetic division... i hope this helps